OpenStudy (gabbyalicorn):

What is the MAD of the data set? {5, 12, 1, 6, 11} I got the first part which is 7 now we have to find the average which i am stuck on help!

OpenStudy (gabbyalicorn):

@Black_ninja123

OpenStudy (gabbyalicorn):

@Here_to_Help15 @geerky42

OpenStudy (gabbyalicorn):

@jabez177

jabez177 (jabez177):

Average, you just add all of them and divide by how many there are.

OpenStudy (gabbyalicorn):

i did that

OpenStudy (gabbyalicorn):

i got 7

OpenStudy (tkhunny):

Step #1 - Find the Average. Step #2 - Find the Absolute Deviation for each number in the data. Just, how far is it from the Mean?

OpenStudy (gabbyalicorn):

but i thought the first step was to do the mean

jabez177 (jabez177):

I'm not that sure. I just got up. Tkhunny can help you on this.

OpenStudy (black_ninja123):

Do you need the second step?

OpenStudy (gabbyalicorn):

Okay, thanks anyway @jabez177

OpenStudy (gabbyalicorn):

yes i do @Black_ninja123

OpenStudy (tkhunny):

It is. Mean = Average If you want the MAD, you need step #2.

jabez177 (jabez177):

You can tag me in a different question. This is a bit too extreme for me.

OpenStudy (gabbyalicorn):

kk

OpenStudy (gabbyalicorn):

but i dont know the second step

OpenStudy (black_ninja123):

Find the absolute value of the differences between each value in the data set and the mean.

OpenStudy (gabbyalicorn):

i got 7 there is no data table or anything :(

OpenStudy (tkhunny):

Yes you do. How far does each entry deviate from the mean? |5-7| = 2 |12-7| = 5 etc... Do all 5.

OpenStudy (black_ninja123):

Use these numbers {5, 12, 1, 6, 11

OpenStudy (gabbyalicorn):

oh so you subtract them with one another.

OpenStudy (black_ninja123):

No subtract them with 7

OpenStudy (tkhunny):

Each one with the Mean. Absolute values only.

OpenStudy (black_ninja123):

5-7 12-7 1-7 6-7 11-7

OpenStudy (gabbyalicorn):

1 - 7 = -6 6- 7 = -1 11 - 7 = 4

OpenStudy (tkhunny):

You seem to be missing the absolute values. This is a MAD, not just a MD.

OpenStudy (black_ninja123):

change the negative sign into a positive sign

OpenStudy (gabbyalicorn):

1 - 7 = 6 ,6- 7 = 1, 11 - 7 = 4

OpenStudy (black_ninja123):

Correct

OpenStudy (gabbyalicorn):

then...?

OpenStudy (black_ninja123):

Find the mean of 2,5,6,1,4

OpenStudy (gabbyalicorn):

3.6

OpenStudy (gabbyalicorn):

is that the answer?

OpenStudy (black_ninja123):

Yes

OpenStudy (black_ninja123):

One sec

OpenStudy (gabbyalicorn):

:D Oh, Yippie!!! Thank you!

OpenStudy (gabbyalicorn):

While you are on AFK @Black_ninja123 ill be working on the next problem. Thanks for your help! ^u^

OpenStudy (black_ninja123):

No problem If you need help just tag me

OpenStudy (gabbyalicorn):

kk ty!